# Optimal Partial Graph Matching

### v2

@@ Appendix B Additional Details on Experiments
As discussed in  Section 8, all the models considered for analysis shown in  Table 2 (right) uses the same neural architecture  (Jiang et al., 2022) to obtain the cross-graph node-to-node affinity matrix and the use Sinkhorn normalization to obtain the doubly stochastic affinity matrix. Once these matrices are obtained, our model OPGM-rs solves the partial graph matching problem as described in  Section 6 by solving a linear sum assignment problem. We use the linear sum assignment problem solver provided by SciPy‘(Virtanen et al., 2020) in our implementation. GCAN (Jiang et al., 2022) solves partial graph matching as an ILP problem, which is usually computationally expensive than solving a LAP.
In the experimental setting we consider (Wang et al., 2023a), OR-tools, which is the Google’s open source software suite for combinatorial optimization has been used to solve the ILP problem. On the other hand, once obtaining the doubly stochastic affinity matrix, AFAT-I and AFAT-I modules use seperate neural modules to first obtain the number of matchings that exist between the two graphs that includes solving an entropic optimal transport problem. Finally, GreedyTopK algorithm (Wang et al., 2023a) which is based on Hungarian algorithm, is used to obtain the final matching. The additional complexities added by separate modules has resulted in higher average inference time of AFAT-I and AFAT-U models, when compared to OPGM-rs.

### v4

@@ Appendix B Additional Details on Experiments
As discussed in  Section 8, all the models considered for analysis shown in  Table 2 (right) uses the same neural architecture  (Jiang et al., 2022b) to obtain the cross-graph node-to-node affinity matrix and the use Sinkhorn normalization to obtain the doubly stochastic affinity matrix. Once these matrices are obtained, our model OPGM-rs solves the partial graph matching problem as described in  Section 6 by solving a linear sum assignment problem. We use the linear sum assignment problem solver provided by SciPy‘(Virtanen et al., 2020) in our implementation. GCAN (Jiang et al., 2022b) solves partial graph matching as an ILP problem, which is usually computationally expensive than solving a LAP.
In the experimental setting we consider (Wang et al., 2023), OR-tools, which is the Google’s open source software suite for combinatorial optimization has been used to solve the ILP problem. On the other hand, once obtaining the doubly stochastic affinity matrix, AFAT-I and AFAT-I modules use seperate neural modules to first obtain the number of matchings that exist between the two graphs that includes solving an entropic optimal transport problem. Finally, GreedyTopK algorithm (Wang et al., 2023) which is based on Hungarian algorithm, is used to obtain the final matching. The additional complexities added by separate modules has resulted in higher average inference time of AFAT-I and AFAT-U models, when compared to OPGM-rs.
@@
IMCPT100IMCPT50Spair71KVOC002020202040404040DatasetsError Rate (%)MismatchingPartialityTotal
Figure 3: Failure mode analysis across four different datasets.
Failure mode analysis
The feasibility of matching a pair of nodes $i$ and $j$ is influenced by their cost $C_{ij}$, the matching biases $\alpha_{i}$ and $\beta_{j}$, and the hyperparameter $\rho$. A specific failure mode we observed is when node pairs that should be matched according to the ground truth are deemed infeasible by the optimization objective in Eq. (9) (as described in Theorem 5.1). Consequently, these nodes remain unmatched in the final solution. This failure mode typically arises when the graph matching network $\textsc{NN}_{\theta}$ assigns an excessively low affinity (resulting in a high matching cost) to pairs of nodes that should otherwise be matched.
We quantify the error by analyzing:
  • Partiality error: The percentage of true matches (as per the ground truth) that fail to meet the feasibility condition outlined in Theorem 5.1. This error reflects mismatches caused by incorrect partiality identification involving matching biases.
  • Mismatching error: The percentage of true matches that are not identified even though they meet the feasibility condition.
Figure 3 demonstrates the partiality error, mismatching error and total error related to this failure mode across four different datasets. The total error is equal to the summation of partiality error and the mismatching error.